a point tranversed half the distace with a velocity a (v0).the remaining part of the distance was covered with velocity v1 for half the time,&with velocity v2for the other half of the time.Find the mean velocity of the point averaged over the whole time of motion.

Grade:

2 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Kaustav, 

let d be the total distance. first half is covered in time d/2v0, let 2t be time for second half, v1t + v2t = d/2

t=d/(2(v1+v2)). mean velocity=d/((d/2v0)+ d/(v1+v2))=2v0(v1+v2)/(v1+v2+2v0)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively :

SAGAR SINGH - IIT DELHI
13 years ago

Dear kaustav,

let D be the total distance. first half is covered in time D/2v0, let 2t be time for second half, v1t + v2t = D/2

t=D/(2(v1+v2)). mean velocity=D/(D/2v0)+ D/(v1+v2))=2v0(v1+v2)/(v1+v2+2v0)

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi